# You Only Sample Once: Taming One-Step Text-to-Image Synthesis by Self-Cooperative Diffusion GANs

### v4

@@ 6 Experiments
In this section, we evaluate the performance of the proposed YOSO. In Sec. 6.1, we evaluate YOSO on unconditional image generation by training from scratch, and we conduct ablation studies in Sec. 6.2 to highlight the effectiveness of our algorithm. Sec. 6.3 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- [6] and Stable Diffusion [40]. In Sec. 6.4, we show that the YOSO can be used for several downstream applications.
Model | FID | NFE | Time (s) |
YOSO | 3.82 | 1 | 0.05 |
Ablation: | |||
YOSO w/o consistency loss | 4.41 | 1 | 0.05 |
YOSO w/o LPIPS loss | 4.63 | 1 | 0.05 |
YOSO w/ | 10.85 | 1 | 0.05 |
UFOGen [60] | 45.15 | 1 | 0.05 |
DDGANs [54] | 3.75 | 4 | 0.21 |
CT [49] | 8.70 | 1 | - |
iCT [48] | 2.83 | 1 | - |
DDPM [12] | 3.21 | 1000 | 80.5 |
EDM [19] | 2.04 | 36 | - |
DDIM [47] | 4.67 | 50 | 4.01 |
StyleGAN2 w/o ADA | 8.32 | 1 | 0.04 |
StyleGAN2 w/ADA | 2.92 | 1 | 0.04 |
6.1 Unconditional Image Generation
Experiment Setting. We use the same UNet and discriminator architecture as used in DDGANs [54]. We set the denoising steps as four. For the evaluation metric, we choose the Fréchet inception distance (FID) [10]. We train the proposed model on the CIFAR-10 dataset [63]. Following Consistency Models [49], we employ LPIPS loss instead of MSE to compute image distance.
Results. The quantitative results are shown in Tab. 1. We observe that our method provides competitive performance compared with the best DMs and GANs. Although some variants of DMs achieve better results, they generally require much longer sample time compared to ours. We also find that among variants of GANs, our performance is surpassed only marginally by StyleGAN2 employing adaptive data augmentation. We also observe that the consistency models provide competitive and even better results compared to us. However, we note that the training of consistency models requires tons of computational resources and careful design of the noise schedule. As shown in Fig. 4, the consistency models require much more computational resources than ours, including more parameters, larger batch size, and more training iterations. To ensure a fair comparison, we re-produce the consistency models by using the same batch size and training iterations as ours. Based on the fair setting, we find that the performance of consistency models significantly decrease, showing way worse performance compared to YOSO. We highlight that the low demand for computational resources enables us to scale the proposed YOSO to large-scale text-to-image datasets.
@@ 6.2 Ablation Studies
We provide additional insight into the effectiveness of the proposed YOSO by performing ablation studies to some potential variants on CIFAR-10 dataset.
@@
Effect of Consistency Loss and LPIPS Loss. We evaluate the effect of consistency loss and LPIPS loss in YOSO. We find that both consistency loss and LPIPS loss can improve the image quality, as indicated by FID (see Tab. 1). However, we notice that if removing consistency loss or LPIPS loss, the generation performance would not decrease significantly. This indicates the effectiveness of our proposed adversarial formulation.
@@
Adversarial Divergence . The key design of our proposed YOSO is smoothing the distribution by using self-generated data instead of real data as ground truth to perform adversarial divergence. We evaluate the variant of using as adversarial divergence. As shown in Tab. 1, the performance of the variant is significantly worse than YOSO. Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant. As shown in Fig. 4, the discriminator loss of the variant is not stable and is much smaller than ours. This indicates that our approach can effectively smooth the distribution to make the distributions of fake and real samples closer to each other, making discrimination harder and hence resulting in higher discriminator loss. Furthermore, we observe mode collapse in the training of the variant. We note that a similar case also happened in SDXL-Turbo, a strong one-step text-to-image model, which also performs adversarial divergence directly against real data (See Fig. 5). These results indicate the superiority of the proposed YOSO.
@@
Comparison to UFOGen. Similar to us, UFOGen does not directly perform adversarial divergence against real data either. However, they perform adversarial divergence over corrupted data (i.e., ). We argue that the formulation is less effective in constructing one-step generation model since their adversarial divergence only indirectly matches the one-step generation at clean data. We implement the UFOGen on CIFAR-10 with our own. Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO. As shown in Tab. 1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID. This further demonstrates the superiority of our proposed YOSO. Through comparison with different adversarial divergence formulations, we find that our formulated adversarial divergence not only smooths the distribution to stabilize the training but also provides effective learning on one-step generation.
6.3 Text-to-Image Generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters. We note that 0.6B parameters can be demonstrated as lightweight among text-to-image DMs. For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec. 5.1. We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Sec. 5.1, followed by training on the JourneyDB dataset [35] with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained.
Evaluation. For image quality, we employ Aesthetic Score (AeS) [45] to evaluate. We also adopt the Human Preference Score (HPS) v2.1 [53] to evaluate the image-text alignment and human preference. The AeS is computed by an aesthetic score predictor trained on LAION [45] datasets, without considering the image-text alignment. HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic. We mainly compare our model against the open-source state-of-the-art (SOTA) models, i.e., SD-Turbo [44] and SDXL-Turbo [44].
Quantitative Results. The quantitative results are presented in Tab. 3. To provide a direct comparison with SD-Turbo, we also fine-tune a model based on pre-trained SD 1.5 on JourneyDB with the introduced technique. We would like to highlight that it is better to conduct training on LAION dataset [45] which is the training set used by SD 1.5. However, it is not available when we conducted this research. Fine-tuning SD 1.5 on other datasets suffers from distribution shift, facing a risk of catastrophic forgetting. As shown in Tab. 3, both YOSO-SD and YOSO-PixArt- clearly beat the SD-Turbo which is based on SD 2.1, a better SD compared to SD 1.5. It is also observed that YOSO-PixArt- outperforms SDXL-Turbo which is a much larger model with 2.6B parameters.
Qualitative Comparison. To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Fig. 5. The comparison shows that: YOSO clearly beats SD-Turbo in terms of both image quality and prompt alignment. Specifically, the samples by SD-Turbo are blurry to some extent. Although SDXL-Turbo with a larger model size improves the sample quality compared to SD-Turbo, we find that the samples by YOSO are sharper than SDXL-Turbo to some extent, and better in prompt alignment. For example, in the first row of Fig. 5 related to a Pikachu, both SD-Turbo and SDXL-Turbo cannot follow the prompt to generate angry expressions and red eyes, while YOSO follows the prompt more precisely. Moreover, it is important to highlight that SDXL-Turbo seems to have a serious issue in mode collapse. Specifically, the third row of Fig. 5 shows that samples by SDXL-Turbo are really close to each other. Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the SOTA one-step text-to-image models.
6.3.1 Zero-Shot One-Step 1024 Resolution Generation
Due to the computational resource limitation, we trained YOSO on 512 resolution. However, there is a 1024-resolution version of PixArt-, obtained by continuing training on the 512-resolution version of PixArt-. Motivated by the effectiveness of the LoRA combination [30], we suggest constructing a similar combination as follows:
6.3.2 Fine-tuning Stable Diffusion via LoRA
We further conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of proposed YOSO. As we perform LoRA fine-tuning, it is expected that it will be worse than full fine-tuning, we propose to use SDXL-Turbo to generate data for LoRA fine-tuning, which avoids the risk of distribution shift to some extent and is economically efficient. We call the model fine-tuned in this way as YOSO-LoRA. We only generate one image for each caption, avoiding the risk of mode collapse brought by SDXL-Turbo. We train YOSO-LoRA with only 64 batch size and 4000 iterations. We present the quantitative results in Tab. 4 and qualitative results in Fig. 6. It can be observed that YOSO-LoRA achieves even better results than LCM with 4 steps and is comparable to InstaFLow with full fine-tuning. And from Fig. 6, YOSO-LoRA has a clearly better sample quality than LCM with 2 steps and InstaFlow. These results indicate that our algorithm enables one-step text-to-image synthesis with reasonable quality by LoRA fine-tuning.
@@ 6.4 Application
One promising and attractive property of DMs is that they can be used in multiple downstream tasks. In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of DMs. Since the relevant applications are mostly developed for SD, we mainly show the results obtained by YOSO-LoRA.
Image-to-image. As shown in Fig. 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing [32] in one step.
@@
Compatibility with ControlNet and Different Base Models. As shown in Fig. 6, our YOSO-LoRA is compatible with ControlNet [64], following the condition well. Our YOSO-LoRA is also compatible with different base models (e.g., dreamshaper and realistic vision) fine-tuned from SD 1.5, preserving their style well. We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step. This may be due to the capability of LoRA and the distribution shift of training data.
Model | NFE | HPS | AeS |
SD 2.1-Turbo [44] | 1 | 27.06 | 5.83 |
Realistic-vision | 50 | 29.76 | 6.35 |
w/ YOSO-LoRA | 4 | 29.50 | 6.21 |
w/ YOSO-LoRA-v2 | 1 | 27.13 | 6.12 |
w/ YOSO-LoRA-v2 | 2 | 29.41 | 6.33 |
Dreamshaper | 1 | 26.64 | 6.08 |
w/ YOSO-LoRA-v2 |
Enabling Data Flywheel. YOSO-LoRA trained on the data generated by SDXL-Turbo, fails to generate reasonable samples in one step when compatible with new base models. In light of the fact that combining YOSO-LoRA with new base models can generate high-quality samples in 4 steps with minimal performance loss as well as preserved style, we propose using YOSO-generated data for training YOSO-LoRA based on existing community models (e.g., Realistic-vision). We call the trained model YOSO-LoRA-v2. As shown in Tab. 5, YOSO-LoRA-v2 can even outperform the full fine-tuned SD-2.1-Turbo in one step. This shows that the YOSO can be trained on slightly shifted data first, and then applied to generate high-quality un-shifted data to train a better YOSO. Moreover, the YOSO-LoRA-v2 can be compatible with some unseen base models (e.g., Dreamshaper) for one-step generation (see Figs. 6 and 5). This shows that combining YOSO-LoRA with community models can generate unshifted training data, enabling the YOSO-LoRA-v2 to focus on learning the acceleration with better compatibility.
@@ Appendix B Limitation
Appendix C Safeguards
The YOSO-SD and YOSO-PixArt- are trained on the JourneyDB dataset [35], which has undergone rigorous human and machine-based filtering to ensure that there are no harmful or violent images in the dataset. YOSO-LoRA and YOSO-LoRA-V2, on the other hand, were trained using data generated by SDXL-turbo and YOSO-LoRA, respectively. To reduce the risk associated with the models, we utilized the built-in safety checker from diffusers [51] to filter out harmful data.

### v5

@@ 6 Experiments
In this section, we evaluate the performance of the proposed YOSO. Sec. 6.1 examines YOSO in the context of text-to-image generation by fine-tuning pre-trained PixArt- (Chen et al., 2024) and Stable Diffusion (Rombach et al., 2022). In Sec. 6.3, we show that the YOSO can be used for several downstream applications. Moreover, we conduct ablation studies in Sec. 6.2 to highlight the effectiveness of our proposed algorithm and proposed training techniques.
6.1 Text-to-Image Generation
Experiment Setting. We initialize the GAN generator by pre-trained PixArt- which is a diffusion transformer with 0.6B parameters. For the GAN discriminator, we construct the latent discriminator by pre-trained SD 1.5 using the proposed construction introduced in Sec. 5.1. We switch the pre-trained PixArt- to v-prediction by the proposed technique introduced in Sec. 5.1, followed by training on the JourneyDB dataset (Pan et al., 2023) with resizing as 512 resolution. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We observe the training convergence fast, requiring only 30k iterations and around 10 A800 days to be trained. We apply a batch-size of 256 and a constant learning rate of 2e-5 during training. We also conduct experiments on fine-tuning SD 1.5 via LoRA to show the effectiveness of the proposed YOSO, with a batch-size of 128 and a constant learning rate of 2e-5.
Evaluation. We employ Aesthetic Score (AeS) (Schuhmann et al., 2022) to evaluate image quality and adopt the Human Preference Score (HPS) v2.1 (Wu et al., 2023) to evaluate the image-text alignment and human preference. The AeS is computed by an aesthetic score predictor trained on LAION (Schuhmann et al., 2022) datasets, without considering the image-text alignment. HPS is trained to predict human preference given image-text pairs, considering the image-text alignment and human aesthetic. Additionally, we include ImageReward score (Xu et al., 2024), and CLIP score (Hessel et al., 2021) to provide a more comprehensive evaluation of the model performance. We mainly compare our model against the open-source state-of-the-art (SOTA) models, e.g., SD-Turbo (Sauer et al., 2023c), PixArt-DMD (Chen et al., 2024; Yin et al., 2023), and Hyper-SD (a concurrent work) (Ren et al., 2024).
Model | Backbone | HFL | Steps | HPS | AeS | IR | CS |
Base Model | SD 2.1 | No | 25 | 27.28 | 5.66 | 0.36 | 33.46 |
SD Turbo (Sauer et al., 2023c) | SD 2.1 | No | 1 | 27.06 | 5.31 | 0.40 | 32.21 |
Base Model | SD 1.5 | No | 25 | 24.72 | 5.49 | 0.20 | 31.88 |
InstaFlow (Liu et al., 2023) | SD 1.5 | No | 1 | 26.18 | 5.27 | -0.22 | 30.04 |
Hyper-SD-LoRA (Ren et al., 2024) | SD 1.5 | Yes | 1 | 28.01 | 5.79 | 0.29 | 30.87 |
YOSO-LoRA (Ours) | SD 1.5 | No | 1 | 28.33 | 5.97 | 0.43 | 31.33 |
LCM-LoRA (Luo et al., 2023b) | SD 1.5 | No | 4 | 22.77 | 5.66 | -0.37 | 30.36 |
PeRFlow (Yan et al., 2024) | SD 1.5 | No | 4 | 22.43 | 5.64 | -0.35 | 30.77 |
TCD-LoRA (Zheng et al., 2024) | SD 1.5 | No | 4 | 22.24 | 5.45 | -0.15 | 30.62 |
Hyper-SD-LoRA (Ren et al., 2024) | SD 1.5 | Yes | 4 | 30.24 | 5.55 | 0.53 | 31.07 |
YOSO-LoRA (Ours) | SD 1.5 | No | 4 | 30.50 | 6.05 | 0.56 | 31.52 |
Base Model | PixArt- | No | 25 | 31.07 | 6.12 | 1.05 | 33.81 |
DMD (Yin et al., 2023) | PixArt- | No | 1 | 29.78 | 6.02 | 0.97 | 32.90 |
YOSO (Ours) | PixArt- | No | 1 | 30.52 | 6.19 | 1.02 | 32.95 |
Quantitative Results. The quantitative results are presented in Tab. 2. As shown in Tab. 2, YOSO clearly beat previous state-of-the-art methods across all metrics (including HPS, AeS, Image Reward score, and Clip Score). It is important to highlight that for the SD 1.5 backbone, we only use LoRA for fine-tuning SD 1.5, where only less than 10% parameters are tuned. Moroever, our method without human feedback learning (HFL) even outperforms Hyper-SD with HFL which potentially hacks the machine metric.
Qualitative Comparison. To provide more comprehensive insight in understanding the performance of our YOSO, we further provide the qualitative comparison in Fig. 4. The comparison shows that: YOSO clearly beats existing baselines in terms of both image quality and prompt alignment, which is achieved by LoRA-finetuning. Specifically, the samples by SD-Turbo are blurry to some extent, and the samples by Hyper-SD are over-saturated and with notable artifacts. Moreover, the advantage of YOSO in text-image alignment is significant. For example, in the last row of Fig. 4 related to a cow-people worker is coding, previous methods including SD 1.5 itself cannot produce samples following the prompt, while YOSO precisely follows the prompt to generate a half cow half people worker. Additionally, it is important to highlight that both SD-Turbo and DMD seem to have a serious issue in mode collapse. As shown in Fig. 5, both SD-Turbo and DMD generate samples that are extremely close to each other. Overall, based on quantitative and qualitative results, we conclude that our YOSO is better in sample quality, prompt alignment, and mode cover compared with the SOTA one-step text-to-image models.
6.1.1 Zero-Shot One-Step 1024 Resolution Generation
Due to the computational resource limitation, we trained YOSO on 512 resolution. However, there is a 1024-resolution version of PixArt-, obtained by continuing training on the 512-resolution version of PixArt-. Motivated by the effectiveness of the LoRA combination (Luo et al., 2023b), we suggest constructing a similar combination as follows:
CM | Lper | IPI | Naive GAN | Coop GAN | DS | AS | HPS | AeS |
✓ | 15.22 | 5.23 | ||||||
✓ | ✓ | 20.08 | 5.45 | |||||
✓ | ✓ | ✓ | 20.63 | 5.57 | ||||
✓ | ✓ | ✓ | ✓ | 24.88 | 5.69 | |||
✓ | ✓ | ✓ | ✓ | 26.42 | 5.85 | |||
✓ | ✓ | ✓ | ✓ | ✓ | 27.10 | 5.94 | ||
✓ | ✓ | ✓ | ✓ | ✓ | ✓ | 28.33 | 5.97 |
@@ 6.2 Ablation Studies
We provide additional insight into the effectiveness of the proposed YOSO by performing ablation studies on some potential variants on CIFAR-10 and on text-to-image generation. We call YOSO without decoupled scheduler and annealing strategy as YOSO-base. See more details in Appendix.
@@
Effect of Consistency Loss. For the effect of consistency loss in YOSO, our results suggest that it can improve the image quality, as indicated by the FID scores in Sec. 6.1.1. Notably, the removal of the consistency loss does not lead to a substantial decline in generation performance, which indicates the effectiveness of our proposed adversarial formulation.
Effect of Latent Perceptual Loss. To investigate the effect of latent perceptual loss proposed in Sec. 5.1, we conduct experiments on text-to-image tasks by using it to compute the consistency loss. As shown in Sec. 6.1.1, the proposed latent perceptual loss clearly improves the HPS from 15.22 to 20.08 and AeS from 5.23 to 5.45, highlighting its effectiveness.
Effect of Decoupled Scheduler. We evaluate the impact of the proposed decoupled scheduler by adding it to YOSO-base. As evidenced by Sec. 6.1.1 and Sec. 6.1.1, the incorporation of the decoupled scheduler significantly enhances sample quality, as reflected by improvements in FID, HPS, and AeS metrics. Notably, the FID score on CIFAR-10 decreases from 3.82 to 3.30 and the HPS in text-to-image tasks is improved from 26.42 to 27.10, demonstrating the superior efficacy of the decoupled scheduler.
Effect of Annealing Strategy. We also assess the effect of the proposed annealing strategy by removing it from the final formulated YOSO. As illustrated in Sec. 6.1.1 and Sec. 6.1.1, the removal of the annealing strategy degrades FID from 3.05 to 3.30 and reduces HPS from 28.33 to 27.10. Such results demonstrate the effectiveness of the annealing strategy in training YOSO.
@@
Adversarial Divergence . The key design of our proposed YOSO is smoothing the adversarial divergence by using self-generated data instead of real data as ground truth to perform adversarial divergence.We evaluate the variant of using as adversarial divergence. As shown in Sec. 6.1.1, the performance of the variant is significantly worse than YOSO-Base, i.e., FID degrades from 3.82 to 10.85. Moreover, besides sample quality, we find that the training of our proposed YOSO is much more stable than the variant. As shown in Fig. 7, the discriminator loss of the variant is not stable and is much smaller than ours. We also compare the variant in the text-to-image task. The results are given in Sec. 6.1.1, from which we can see that replacing our cooperative adversarial loss with yields serious performance deterioration, i.e., HPS decreases from 26.42 to 24.88 and AeS decreases from 5.85 to 5.69. This indicates that our approach can effectively smooth the adversarial divergence by narrowing the gap between fake and real distribution, which makes discrimination harder and hence increases discriminator loss. Furthermore, we observe mode collapse in the training of this variant. Note that a similar case also occurs in SD-Turbo, a strong one-step text-to-image model, which performs adversarial divergence directly against real data (See Fig. 4). Overall, these results indicate the superiority of the proposed YOSO.
@@
Comparison to UFOGen. Similar to us, UFOGen does not directly perform adversarial divergence against real data either. Instead, they employ adversarial divergence over corrupted data (i.e., ). However, we argue that such a formulation is less effective in constructing one-step generation model as the adversarial divergence it employs only indirectly aligns with one-step generation at the level of clean data. We implement UFOGen on CIFAR-10 on our own. Note that we also perform consistency loss in the UFOGen variant to ensure a fair comparison, as the consistency loss has been found useful in YOSO. As shown in Sec. 6.1.1, UFOGen only obtains 45.15 FID on CIFAR-10, significantly worse than YOSO with 3.82 FID. This further underscores the superiority of our proposed YOSO. Through comparison with different adversarial divergence formulations, we find that our approach not only stabilizes training by smoothing the adversarial divergence but also facilitates effective learning for one-step generation.
@@ 6.3 Application
One promising and attractive property of DMs is that they can be used in multiple downstream tasks. In this section, we show the capability of our proposed YOSO in various downstream tasks, keeping the unique advantages of DMs: 1) Image-to-image Editing: As shown in Fig. 6, our YOSO-LoRA is capable of performing high-quality image-to-image editing (Meng et al., 2022) in one step; 2) Compatibility with ControlNet and Different Base Models: As shown in Fig. 6, our YOSO-LoRA is compatible with ControlNet (Zhang et al., 2023), following the condition well. Our YOSO-LoRA is also compatible with different base models (e.g., Dreamshaper and Toonyou) fine-tuned from SD 1.5, preserving their style well. We find that when applying YOSO-LoRA to new base models, it fails to produce reasonable samples in one step. This may be due to the capability of LoRA and the distribution shift of training data.
7 Related Works
Few-Step Text-to-Image Generation. LCM (Luo et al., 2023a) adapts consistency distillation (Song et al., 2023) for stable diffusion (SD) and enables image generation in 4 steps with acceptable quality. InstaFlow (Liu et al., 2023) adapts rectified flows (Liu et al., 2022; Liu, 2022) for SD, facilitating the generation of text-to-image in a mere single step. Despite this, the fidelity of the generated images is still poor. Some works (Yin et al., 2023) adapts variation score distillation (Wang et al., 2023) and diff-instruct (Luo et al., 2023c) for SD, which requires training an additional DMs for the generator distribution, akin to the discriminator in GANs. We found models trained in this way seem to have serious mode collapse issues (see Fig. 5). Recently, combining DMs and GANs for one-step text-to-image generation has been explored. UFOGen (Xu et al., 2023c) extends DDGANs (Xiao et al., 2022) and SSIDMs (Xu et al., 2023b) to SD by modifying the computation of reconstruction loss from corrupted samples to clean samples. However, it still performs adversarial matching using corrupted samples. ADD (Sauer et al., 2023c) introduces a one-step text-to-image generation based on SD. It follows earlier research (Sauer et al., 2023a) by employing a pre-trained image encoder, DINOv2 (Oquab et al., 2024), as the backbone of the discriminator to accelerate the training. However, the discriminator design moves the training from latent space to pixel space, which substantially heightens computational demands. Moreover, they directly perform adversarial matches at clean real data, increasing the challenge of training. This requires the need for a better but expensive discriminator design and expensive R1 regularization to stabilize the training. SDXL-Lightning (Lin et al., 2024b) combines progressive distillation and adversarial training in a progressive framework. Recently, a concurrent work Hyper-SD (Ren et al., 2024) combines consistency distillation and adversarial training in a progressive framework, while our work can be trained end-to-end. However, their work performs adversarial point-level matching via consistency loss, and they still perform adversarial training via injecting noise. This results in less effective one-step learning, thus they rely on human feedback learning for achieving better performance. In contrast, we perform the adversarial training at the distribution level by replacing real data with self-generated data to smooth the adversarial divergence, which not only can stabilize the training without injecting noise but also form effective one-step generation learning. Moreover, compared with above mentioned diffusion-GAN hybrid models, our approaches can be trained from scratch to perform one-step generation, which is not demonstrated by them. Additionally, we extend our method not only to SD but also to PixArt- (Chen et al., 2024) which is based on diffusion transformer (Peebles & Xie, 2022). This demonstrates the wide application of our proposed YOSO.
@@ Appendix B Limitation
Appendix C Proof
Proof: The cooperative adversaril loss is defined as: . Since the divergence is non-negative, and the is defined as , there exist a global optimal solution such that = 0. This completes the proof.